OpenStudy (anonymous):

How do you do directed numbers:) Place the correct symbol < or > between each pairs of numbers +7 something _9

OpenStudy (anonymous):

plzz help

OpenStudy (amistre64):

the bigger end is towards the bigger number the smaller end is towards the smaller number

OpenStudy (amistre64):

so, the question is: which is bigger, +7 or -9?

OpenStudy (anonymous):

so this is < big and dis > is small

OpenStudy (amistre64):

lets me give an example: 10 is bigger than 3 10 > 3 is correct 3 < 10 is correct each time the small end is at the small number, and the large end is at the larger number

OpenStudy (anonymous):

Yes I get it now:) Thanks so much! Can i ask u another question? please

OpenStudy (amistre64):

one more i spose

OpenStudy (anonymous):

place each set of numbers. start with the smallest +9, -5, -8

OpenStudy (amistre64):

it might help to think of this as addition and subtraction; think of a number that is larger than 9, 5, or 8 ; say 10 now we can view this as: 10 + 9 = 19 10 - 5 = 5 10 - 8 = 2 the order from smallest to largest of these will match the order of smallest to largest of your original question

OpenStudy (amistre64):

2 < 5 < 19 ; just replace each of these with the value we modified "10" by to get them -8 < -5 < 9

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome